### v1

# Entity-based Reinforcement Learning for Autonomous Cyber Defence

Abstract
A significant challenge for autonomous cyber defence is ensuring a defensive agent’s ability to generalise across diverse network topologies and configurations. This capability is necessary for agents to remain effective when deployed in dynamically changing environments, such as an enterprise network where devices may frequently join and leave. Standard approaches to deep reinforcement learning, where policies are parameterised using a fixed-input multi-layer perceptron (MLP) expect fixed-size observation and action spaces. In autonomous cyber defence, this makes it hard to develop agents that generalise to environments with network topologies different from those trained on, as the number of nodes affects the natural size of the observation and action spaces. To overcome this limitation, we reframe the problem of autonomous network defence using entity-based reinforcement learning, where the observation and action space of an agent are decomposed into a collection of discrete entities. This framework enables the use of policy parameterisations specialised in compositional generalisation. Namely, we train a Transformer-based policy on the Yawning Titan cyber-security simulation environment and test its generalisation capabilities across various network topologies. We demonstrate that this approach significantly outperforms an MLP-based policy on fixed networks, and has the ability for zero-shot generalisation to networks of a different size to those seen in training. These findings highlight the potential for entity-based reinforcement learning to advance the field of autonomous cyber defence by providing more generalisable policies capable of handling variations in real-world network environments.
1 Introduction
The development of reliable agents for autonomous cyber defence is a growing research area aimed at addressing the increasing complexity and volume of cyber threats. Solutions based on deep reinforcement learning (RL), inspired by high-profile successes in other complex domains (Vinyals et al.,, 2019; Roy et al.,, 2021; Bello et al.,, 2017), offer the potential to rapidly detect, analyse, and respond to cyber attacks at scale (Nguyen and Reddi,, 2021). A fundamental challenge involved in the deep RL approach is to develop agents that are generalisable to variable network topologies. This is difficult to achieve with fixed-input neural network architectures commonly used for function approximation and standard RL interfaces such as Gym or Gymnasium (Brockman et al.,, 2016; Towers et al.,, 2024) which expect fixed-size observation vectors and action spaces. A lack of this generalisation ability imposes serious limitations on the potential deployment of RL agents in real-world network scenarios, as enterprise networks can vary dramatically in size, structure, and complexity.
One way of approaching RL for autonomous cyber defence is to construct an observation vector by concatenating features from all of the nodes in the network, and applying standard deep RL algorithms with policies or value functions parameterised using multi-layer perceptrons (MLP). Such an agent can be trained in environments with fixed-size observation and action spaces (Standen et al., 2022a, ), but would not be readily transferable to other configurations. For example, under this approach, it would be impossible to deploy an agent on networks larger than the ones encountered during training, and any attempts to adapt to different networks that are the same size or smaller using padding are unlikely to be successful due to the ‘non-exchangeability’ of input features  (Mern et al.,, 2020, 2019). That is, the nature of nodes and connections between them also affect the meaning of elements of the agent’s observations, with elements in the same position in the input vector to the policy network not necessarily representing a single consistent object. Also, actions indexed in a discrete action space may not have a consistent effect depending on the objects they interact with. This places a large burden on an MLP to learn these interactions from scratch, which may require an impractically large number of environment samples.
This problem falls under the broad category of compositional generalisation, which is a long-studied problem in reinforcement learning (Boutilier et al.,, 2000; Guestrin et al.,, 2003; Diuk et al.,, 2008). Recently, a number of works have explored introducing compositional generalisation to deep reinforcement learning  (Lin et al., 2023a, ; Mambelli et al.,, 2022; Zhou et al.,, 2022; Haramati et al.,, 2024), through using permutation invariant policy architectures based on attention (Vaswani et al.,, 2017) or Deep Sets (Zaheer et al.,, 2017). Indeed, some existing works relating to autonomous cyber defence have suggested using Transformers to parameterise a value function (Mern et al.,, 2021), or using graph neural networks for policies (Janisch et al., 2023a, ; Janisch et al., 2023b, ; Palmer et al.,, 2023) which are invariant to the number of nodes in the environment.
This work aims to address this challenge in the context of autonomous cyber defence by leveraging entity-based RL (Winter et al., 2023a, ). Entity-based RL is a paradigm where environments are decomposed into collections of discrete entities which an agent observes and acts upon directly. In the context of a network environment, each node of the network can be treated as an entity, with a defending agent’s observation space permitted to vary between environment instances according to how many nodes are visible. With a suitable training regime, it should be possible for an agent trained using entity-based RL to generalise to any network topology, provided the network in question contains familiar types of entities. The Entity-Gym framework (Winter et al., 2023b, ), intended to occupy a similar role as OpenAI Gym or Gymnasium, provides a standard interface in which to cast entity-based environments, and allows any agent designed for entity-based reinforcement learning to be seamlessly deployed on such environments.
In this paper, we introduce a wrapper for the Yawning Titan cyber simulation environment (Andrew et al.,, 2022) that adheres to the Entity Gym interface. On this environment, we trained and compared two types of policies: a Transformer-based policy, RogueNet (Winter et al., 2023a, ), trained on the entity-based interface and a multilayer perceptron (MLP) policy trained on the standard Gym (Brockman et al.,, 2016) interface for Yawning Titan. We trained both types of policies using Proximal Policy Optimization (PPO) (Schulman et al.,, 2017) with two different training modes, which we refer to as Random and Static, across network sizes of 10, 20, and 40 nodes. In the Random mode, a new random network topology is generated with each environment reset, whilst in the Static mode, a random network topology is initialised at the start of training and kept constant throughout. This approach has been designed to facilitate comparison with the MLP-based policy, which is able to train well in the Static mode, but struggles with Random training where it encounters a new network configuration with every episode. We also performed zero-shot evaluation of the Transformer-based policies on network sizes not seen during training, something not possible with standard MLP-based policies. From these experiments we find that in this environment, entity-based policies outperform MLP-based policies during training, particularly in the Random mode where the MLP-based policy struggled to learn. We also find that entity-based policies generalise well to networks of unseen sizes, performing similarly to policies trained natively in those environments.
The paper is organised as follows. In section 2 we give an overview of motivating background and related literature. In section 3 we describe the Entity Gym and the RogueNet architecture (Winter et al., 2023a, ; Winter et al., 2023b, ) which we have used to construct our environment and train entity-based policies. In section 4 we provide a description of the Yawning Titan simulator, and choices we have made for the environment setting for our experiments. In section 5 we explain how we have adapted the Yawning Titan simulator to the entity-based framework. In section 6 we outline the methodology and experiments we have chosen to run. In section 7, we present the results of our experiments and conclude with a discussion of results and future work in section 8.
2 Background and related work
2.1 Autonomous Cyber Defence
Over the last decade, a number of simulators have been developed to facilitate research into using reinforcement learning for autonomous cyber defence (Standen et al., 2021a, ; Standen et al., 2022a, ; TTCP CAGE Working Group,, 2023; Standen et al., 2021b, ; Molina-Markham et al.,, 2021; Andrew et al.,, 2022; Microsoft Defender Research Team,, 2021). Typically, these involve a network scenario consisting of nodes of varying types. A defensive blue agent is provided with a fixed observation and action space with which to learn to defend the network from an attacking red agent, whose behaviour is unintelligent and part of the environment dynamics.
The CybORG environment (Standen et al., 2022b, ) was initially conceived as part of a competition to encourage research into RL-based autonomous cyber defence agents, Cage Challenge 1 (Standen et al., 2021a, ). High scoring solutions to the challenge used agents trained with PPO (Schulman et al.,, 2017), combined with a bandit controller to select between policies specialised in specific attack patterns (Foley et al., 2022a, ; Hicks et al.,, 2023; Foley et al., 2022b, ; Bates et al.,, 2023), or with an ensemble approach (Wolk et al.,, 2022). Whilst not the focus of our paper, the CybORG environment features several types of nodes that could be used as entity types in the entity-based RL setting; User Hosts, Enterprise Servers, Operational Hosts and the Operational Server.
Recognising that training and evaluating on a purely fixed network environment is not satisfactory for real-world applicability, a number of works have explored forms of generalisation. For example, to unseen red agent dynamics (Wolk et al.,, 2022), or to unseen network topologies, whether that be on fixed networks with variable connectivity (Collyer et al.,, 2022), or to networks containing differing numbers of nodes (Applebaum et al.,, 2022; Mern et al.,, 2021; Janisch et al., 2023a, ; Janisch et al., 2023b, ; Palmer et al.,, 2023). Indeed, the recent CAGE Challenge 4 (TTCP CAGE Working Group,, 2023) introduced a level of variability in the number of possible user nodes to encourage solutions addressing this problem.
One suggestion (Palmer et al.,, 2023) is to use graph neural-networks (GNNs), specifically Graph Attention Networks (Velickovic et al.,, 2018) (GATs) to exploit the structural information available in cyber environments as well as imbue any defensive policy with some degree of invariance to the number of nodes in a network. For penetration testing using an autonomous network attacker, Janisch et al., 2023a use an ‘invariant’ architecture for the attacking agent’s policy. The agent has a shared encoder for every node, and amalgamates information from nodes as they are discovered by the agent. This is similar to the message passing seen in GNN-based policies. In Janisch et al., 2023b , the same authors study the use of a GNN-based policy, together with autoregressive policy decomposition for the SysAdmin game (Guestrin et al.,, 2003), and test the generalisation ability of the policy on networks of variable size.
To the best of our knowledge, the most relevant prior work is Mern et al., (2021, 2022), where the authors implement a reinforcement learning environment for the autonomous defence of an industrial control network, and use an attention-based action-value function. Similarly to the ‘entity-based’ paradigm, they use separate modules to encode different node-types, and share encoder parameters between all nodes of the same type. A global attention layer is then used to pass information between all node embeddings, before action-value estimates are decoded for all actions on all nodes. The parameters are learnt using a version of the Deep Q-Learning algorithm (Mnih et al.,, 2013; Hessel et al.,, 2017). This differs from the explicit policy parameterisation used in this work, trained using PPO. An advantage of a policy-based approach is that it allows for a more natural implementation of composite action spaces.
2.2 Entities and Objects in Reinforcement Learning
In this section we provide an overview of attempts to leverage the underlying structure of RL environments, as well as practical approaches for dealing with complex environments with potentially variable observation and action spaces, such as in the case of autonomous cyber defence. In the first instance, particularly in the case of Factored MDPs Boutilier et al., (2000), the motivation is to find more efficient solution methods by exploiting the ‘factorised’ structure of an environment. More broadly, all of these approaches contribute to solving the problem of compositional generalisation (Zhao et al.,, 2022; Lin et al., 2023b, ). That is, training policies that are able to generalise across environments containing different compositions of familiar objects or entities and dynamics. For further reference, Mohan et al., (2024) provide a comprehensive survey of structure in reinforcement learning.
2.2.1 AlphaStar and entity-based Reinforcement learning
AlphaStar (Vinyals et al.,, 2019; Choi,, 2020) is an agent developed to play the game Starcraft II at human professional level. Starcraft II is a real-time strategy game where two opposing players develop a base of operations and direct armies of units with the goal of defeating their opponent. Starcraft is a hard problem, both in terms of the practicalities of agent design and the complexity of strategies, which captures many of the challenges of real-world deployment of autonomous agents. Of particular relevance is that the number and nature of the entities an agent can observe and act upon varies throughout the game. This problem may be similar to that of nodes joining or leaving a network, or more generally the problem of transferring to an unfamiliar network consisting of familiar entities. Rather than using purely image-based inputs of the game screen as observations, AlphaStar maintains an entity list consisting of every visible object the agent can interact with. This could be army units, buildings, resources, or terrain. Crucially, this list will vary throughout a game, and between games.
Each entity is associated with a one-dimensional vector that summarises the state and properties of the entity. AlphaStar encodes these entity observations through a three-layer Transformer (Vaswani et al.,, 2017) to learn the relationships between entities. The entity embeddings generated by the Transformer are merged with other scalar and game-map information to form a joint representation for a single time-step’s observation, which is used as input to an LSTM (Hochreiter and Schmidhuber,, 1997) that learns temporal dependencies.
Since the actions available to the agent depend on the entities that are visible, the policy must be equipped to deal with a variable-size action space, and to map information directly from its entity-based observations to the relevant actions associated with each entity. This is achieved through using skip-connections from the entity embeddings to an entity-selection head. This network computes a key for each entity from the provided entity embeddings, which are then used by a recurrent pointer network (Vinyals et al.,, 2015) to select which entities to act on. This pointer network is invariant to the number of entities available at any given time. Similarly, there is a target entity head which works the same way, but selects a single entity, for actions that might first involve selecting entities and then targeting another. In Starcraft, this could be selecting a group of friendly units to target an opponent.
2.2.2 Factored MDPs
Factored MDPs (Boutilier et al.,, 2000; Guestrin et al.,, 2003; Osband and Van Roy,, 2014) provide a framework that allows for more efficient learning on environments that may be naturally ‘factorised’ into distinct components. Here, the reward and transition functions can also be factorised according to these components, and display conditional independence properties. Guestrin et al., (2003) introduced efficient solution methods for reinforcement learning on factored MDPs as well as a toy environment, SysAdmin, on which to study them. In this environment, a system administrator must maintain a network of computers, each of which can be functional or failed. The administrator must decide whether to reboot one of the machines or do nothing at each time step, aiming to maximise rewards received for working machines, while considering that machine failures are influenced by the state of neighbouring machines. In this environment, the transition of each individual machine is conditionally independent of all other machines given its immediate neighbours, hence it can be cast as a factored MDP. The SysAdmin example instinctively bears resemblance to autonomous cyber defence environments, and F-MDP framework can be a suitable candidate to express network security problems such as the ones considered in this work.
2.2.3 Relational Reinforcement Learning
Relational reinforcement learning (RRL) is the study of solution methods for environments where states and actions can be naturally represented using relational and logical structures (Džeroski et al.,, 2001). This framework aims to address the limitations of traditional RL methods when dealing with structured environments and to enable more efficient generalisation across tasks.
In RRL, the state space, action space, and transition function are represented using relational descriptions. This allows for a more compact and expressive representation of complex environments. The key idea is to learn relational policies that can generalise across objects and relations, rather than learning separate policies for each specific instance van Otterlo, (2005).
RRL has been applied to various domains, including robotics walker2004relational, game playing guestrin2003generalizing, and natural language processing branavan2012learning. Its ability to handle structured environments and transfer knowledge across tasks makes it a promising approach for developing more generalisable and sample-efficient reinforcement learning algorithms.
Zambaldi et al., (2019) study ‘deep’ relational reinforcement learning, and advocate the use of GNN or attention-based function approximators in order to learn abstract representations of relations between entities in an environment. This is opposed to traditional relational RL, where relations are provided a-priori. They find the use of ‘relational’ attention modules allows an agent to generalise zero-shot to unseen environment configurations. This approach essentially discards the traditional formal solution methods based on logical descriptions of the environment. In other words, it is a purely ‘entity-based’ approach, where the environment is split into distinct entities, and any relations between them are learnt during training.
Similarly, Janisch et al., 2023b study the use of a GNN-based policy on relational planning domains, including the ‘SysAdmin’ environment introduced for the study of factored MDPs by Guestrin et al., (2003), as mentioned above. They find their approach demonstrates zero-shot generalisation ability in environments with larger numbers of objects than seen during training, including ‘SysAdmin’ networks of variable sizes. Motivated by this, Nyberg and Johnson, (2024) recently applied this GNN-based policy architecture to the CAGE Challenge 2 cyber defence environment (Standen et al., 2022a, ). The authors experiment with modifying the default scenario by adding and removing user-level nodes, and testing the zero-shot generalisation performance of the GNN-based policy on these variants.
2.2.4 Object-oriented Reinforcement Learning
Similarly to RRL, Object-Oriented Reinforcement Learning (OORL) (Wasser,, 2010) offers a formalism for representing RL environments consisting of distinct objects, where each object belongs to a class with a particular set of features or attributes. That is, objects that fall within a certain class $c_{i}\in\mathcal{C}=\{c_{1},...,c_{m}\}$ possess a particular set of class-specific attributes $\{c_{i}.a_{1},...,c_{i}.a_{k}\}$. It is similar to the RRL formalism, but less restrictive in terms of defining the interactions between objects. The ‘object-oriented’ or ‘object-centric’ RL terminology has been adopted in research into visual robotics tasks. Often where the objects in the environment are discovered using an object detection architecture, before being fed into a GNN or Transformer-based policy (Janner et al.,, 2019; Mosbach et al.,, 2024; Locatello et al.,, 2020; Haramati et al.,, 2024), as opposed to being defined as part of the environment in advance. Similarly to deep RRL, the terminology has survived, but when using solutions based on deep learning with transformers or GNNs, the specifics of the formalism is not particularly relevant. What is consistent is the recognition that partitioning an environment into distinct objects with their own features, and training a permutation invariant set-based architecture across these objects results in more efficient training and generalisation capabilities.
2.2.5 Graph-based Reinforcement Learning for robotic control
A number of papers have studied the use of GNN or Transformer-based policies on robotics continuous control environments, which Hong et al., (2022) refers to as ‘Inhomogeneous Multitask Reinforcement Learning’. Here, the agent is tasked with learning policies across multiple tasks that may have different (inhomogeneous) state and action spaces or dynamics. The goal is to leverage shared knowledge and experience across these tasks to improve learning efficiency and performance. This has been successfully applied to robotics environments (Wang et al.,, 2018; Huang et al.,, 2020; Kurin et al.,, 2021; Hong et al.,, 2022), where the observation and action space are taken to be nodes in a graph of joints of the robot. An agent may be trained across different robot morphologies given by different graphs, with the goal that the policy is generalisable across all of these settings. Such tasks can be formalised by considering a set of MDPs $\mathcal{M}^{K}=\{M_{1},\dots,M_{K}\}$ (Kurin et al.,, 2021; Hong et al.,, 2022). The set $\mathcal{M}^{K}$ is inhomogeneous if there exists $i,j\in[K],i\neq j$ such that $dim(\mathcal{S}_{i})\neq dim(\mathcal{S}_{j})$ or $dim(\mathcal{A}_{i})\neq dim(\mathcal{A}_{j})$, where $\mathcal{S}_{i}$ and $\mathcal{A}_{i}$ refer to the state and action spaces of MDP $M_{i}$. The aim is then to construct a policy that maximises the average expected discounted return over all environments in $\mathcal{M}^{K}$, i.e., $\frac{1}{K}\sum_{i=1}^{K}\mathbb{E}[\mathcal{R}_{i}]$, where $\mathcal{R}_{i}$ denotes expected returns in MDP $M_{i}$. In order to train such policy, Wang et al., (2018) propose to use GNNs to model the relationships between different components of the state and action spaces across tasks, as GNNs are capable of capturing the structural and relational information inherent in various robot morphologies.
2.2.6 Generalisability in Single-Agent RL
Finally, we discuss some relevant popular work on context generalisability in single-agent RL Kirk et al., (2023). Within this domain, contributions such as Zhang et al., (2020); Seitzer et al., (2021) and Han et al., (2021); Ghosh et al., (2021) have focused on developing methods that expose the agent to a broad range of contextual episodic variability during training, enabling it to learn robust and invariant causal representations across different ‘contexts’ (mostly across episodes). These policies are then proven to generalise better to new, unseen contexts at test time. The difference between these methods and the entity-based approach used in this work is mainly conceptual and lies in the distinction between the time-invariant context and the entity-level representations that can vary within-episode, as they are treated as part of the evolving environment’s state space.
3 Entity-based Reinforcement Learning
To apply entity-based RL to Yawning Titan, we make use of the main constituents of the Entity Neural Network project of Winter et al., 2023a . That is, the Entity Gym interface (Winter et al., 2023b, ), the RogueNet policy architecture (Winter et al.,, 2021), and the Entity Neural Network Trainer framework for training agents in Entity Gym environments (Winter et al., 2023c, ). The following section is a summary of components of these works necessary for later discussion and doesn’t represent an original contribution of this paper. For more details on implementation and architecture, refer to the documentation of the repositories of Entity Gym (Winter et al., 2023b, ) and the Entity Neural Network Trainer (Winter et al., 2023c, ), as well as the blog post (Winter et al., 2023a, ) providing a comprehensive introduction to entity-based RL and the motivation behind the Entity Neural Network project.
3.1 Entity Gym
In this section we describe the features of the Entity Gym framework (Winter et al., 2023a, ; Winter et al., 2023b, ) and the choices available for specifying an entity-based environment.
Environment
The Environment class in Entity Gym defines a standard interface for entity-based RL. Following a similar structure to Gymnasium (Towers et al.,, 2024), this interface includes methods for defining the observation and action spaces, resetting the environment, processing actions, and generating observations. To create a custom entity-based environment, one inherits from the Environment class and implements these core abstract methods. This process involves defining all relevant entity types and their features, as well as the action spaces associated with specific entity types. As we do in this paper, if replicating an existing environment such as Yawning Titan which has a Gym or Gymnasium interface, it’s important to ensure that the observations and actions in the Entity Gym interface correspond to those in the original version and that functionality is preserved. In other words, it should be the case that the interface and representation of the environment is different, but the information provided to the agent and the fundamental dynamics of the environment are the same.
Entities
Any entity-based environment is initialised with lists of possible entities, all with a specified type. Each entity type is simply defined by a list of named features that are associated with that entity type. Observations
Each observation can contain global features, entity-specific features, action masks and rewards for the current timestep. The entity features are stored in a dictionary mapping entity types to lists of feature arrays, where each element in the list are the features for a particular entity. These lists are permitted to be any length at any given timestep, allowing for variable numbers of entities in the environment. Action masks, defined for each action type, specify which entities can perform actions and what actions are available at a particular timestep. These masks directly interact with the action space, constraining the set of valid actions for each entity. We don’t make use of action masks for our environment, although one might imagine them being useful for a real cyber environment, where access to particular nodes may be prevented.
Action Spaces
Three kinds of actions spaces are supported:
  • GlobalCategoricalActionSpace’: For global actions that aren’t associated with specific entities.
  • CategoricalActionSpace’: This action space allows the agent to direct a set of entities to execute a discrete action, where a different action may be executed for each entity. Each instantiation of this action space is associated with a particular entity type, or set of entity types.
  • SelectEntityActionSpace’: This action space allows a set of entities to choose another entity to act on. Similarly, each instantiation of this action space is associated with particular actor and actee entity types.
For the GlobalCategoricalActionSpace, one might assign actions that affect a large number of entities simultaneously, or have no direct effect on the entities in the environment but change the state of the agent. For example, moving the top-down camera in a video game such as StarCraft Vinyals et al., (2019), which changes the entities that are visible to the agent. For an autonomous network-based agent, this could be any action that has network-wide implications, or perhaps the decision to intervene in the network if malicious activity is detected Hammar and Stadler, (2022).
For the ‘non-global’ CategoricalActionSpace, one might assign actions that are associated with an entity-level effect. In the example of a grid-world game, this could be moving entities in canonical directions. In an autonomous cyber defence environment, this could be node-level actions such as rebooting specific nodes, or terminating specific processes.
SelectEntity actions could be used for anything that involves directing one entity to interact with another. In AlphaStar, the agent has a “Target unit” head in its action pipeline, which allows collections of entities to apply a previously selected action to another entity. In the cyber-defence setting one might imagine implementing an action that allows a centralised agent to select one node to send keying material, a digital certificate or a backup to another node. Additionally, one might imagine treating services or users on a particular node as entities in themselves, and directing the node to select one of its internal entities to act on (for example, terminating a service). One could also define any defensive agent or node issuing defensive commands as an abstract entity in itself, and treat all defensive actions involving operations on specific nodes as SelectEntityActions, where the defensive entity selects other nodes to perform actions on. This can be useful for constructing complex action spaces, in combination with the Categorical or GlobalCategorical action spaces, where an action type could be chosen separately from the node to act on.
3.2 RogueNet
RogueNet (Winter et al.,, 2021) is a Transformer-based policy network implementation, designed to work specifically with Entity Gym compatible environments (Winter et al., 2023b, ), and uses a specially implemented ragged batch data-type (Winter,, 2023) to deal efficiently with variable length observations as input. It is similar to the entity encoding architecture used by Vinyals et al., (2019) (see 2.2.1).
Input Representation
RogueNet takes ‘entity-based’ observations as input, represented as a set of features for each entity. The input is structured as a mapping from entity types to ragged batch objects, allowing for efficient training with batching whilst maintaining support for variable numbers of entities across environment instances. These batches are constructed from the lists of entity features that constitute the observations provided to the agent at each timestep. Masks are used to prevent information leakage between timesteps and environment instances. Global features, such as network connectivity information, could be incorporated and are appended to each entity’s feature set, maintaining a consistent input structure.
Entity Embedding
An entity embedding head is initialised for every entity type in the environment. The same embedding parameters are used for all entities of a given type. These heads are single-layer MLPs with input and layer normalisation, and ReLU nonlinearity.
Transformer Backbone
After entities are embedded, these embeddings are passed through several layers of Transformer blocks, in our case 2. Each block comprises self-attention and feedforward sublayers with residual connections and layer normalisation.
Action Heads
RogueNet uses multiple action heads based on the action spaces defined in the environment, with the three types defined in section 3.1. Both GlobalCategorical and Categorical actions use the same categorical action head implementation. For global categorical actions, a dummy ‘global’ entity is introduced. This entity’s representation is updated in each Transformer layer along with other entities, allowing it to aggregate information from all entities. This action head is implemented as a simple affine layer followed by a softmax, taking as input the entity-embedding after being fed through the Transformer backbone of the policy. For SelectEntity actions, the action head computes a query $q$ from the embedding of the actor entity, and key $k$ vectors for each selectable actee entity:
$q=W_{q}h_{\text{actor}},\quad k_{i}=W_{k}h_{i}$
where $W_{q}$ and $W_{k}$ are learnable weight matrices, $h_{\text{actor}}$ is the representation of the acting entity and $h_{i}$ are the representations of candidate actee entities. A dot product between these embeddings is then used to create a probability distribution over entities. This is similar in function to pointer networks (Vinyals et al.,, 2015). That is, the probability of selecting entity $i$ is determined by $\frac{q^{T}k_{i}}{\sqrt{d_{qk}}}$ where $d_{qk}$ is the dimension of the query and key vectors, which is passed through a softmax over all eligible entities $i$ to produce valid probabilities.
Training and Inference
RogueNet can be trained using policy gradient methods, and we use the version of PPO implemented in the companion Entity Neural Network Trainer package (Winter et al., 2023c, ), which is itself adapted from CleanRL (Huang et al., 2022b, ). RogueNet’s use of ragged tensors and the attention mechanism enables it to process varying numbers of entities without altering its components or structure. This property allows the architecture to be applied to environments with dynamic entity populations or across multiple environments with different numbers of entities, such as network environments with different numbers of nodes. It is straightforward to deploy a trained policy on any Entity Gym environment containing the same types of entities and action spaces as the environment used for training the policy. For our use-case, this is network environments containing variable numbers of nodes.
4 Yawning Titan
Figure 1: Plots showing examples of the structure of random networks used in the Yawning Titan environment, with entry nodes marked in red.
Yawning Titan is an abstract, graph-based environment designed for the development of RL agents for autonomous cyber defence (Andrew et al.,, 2022). It provides an OpenAI Gym (Brockman et al.,, 2016) interface for a network-based cyber defence game through which a defending agent (blue agent) can interact with a network environment containing an attacker (red agent). Yawning Titan allows for flexibility in the configuration of the network and the dynamics of the game, including types of nodes, network topology, actions and observations available, termination, rewards and red agent behaviour. Several works have used and adapted Yawning Titan to train and evaluate autonomous blue agents (Acuto and Maskell,, 2023; Andrew et al.,, 2022; Collyer et al.,, 2022) with various different environment configurations. In the following, we describe the main components of the environment configuration we have used for the experiments in this paper, following previous work closely (Collyer et al.,, 2022) but with some differences to allow for clearer comparisons. Yawning Titan includes more features and options for configuration than discussed here. More information can be found in the documentation and repository, as well as the original paper (Andrew et al.,, 2022).
4.1 Nodes and Network
The nodes in the environment are simple objects with two main attributes, Node Vulnerability and Compromised Status. Node Vulnerability is a value in the range $[0,1]$ that dictates how easily a node can be compromised by an attacking red agent, with a higher value meaning the node is more vulnerable. Each node might also have an additional special attribute:
  • High Value / Target Node: A flag that triggers special environment behaviour when the node is compromised, such as episode termination or a large negative reward for the blue agent.
  • Entry Node: At the start of an episode, the red agent may only enter the network by attacking designated entry nodes. From there, it can attack nodes adjacent to those it has compromised already.
  • Deceptive Node: It is possible to configure the environment such that the blue agent may add new deceptive nodes to the network, which divert the red agent away from compromising the true network.
For our experiments we used a single randomly selected Entry Node, and we did not use the High Value Node and Deceptive Node functionality. We mention them here as these are examples of possible varieties of entity type that could be included in more complex environments to make full use of the entity-based framework. Adding the deceptive node functionality would also mean that the number of nodes in the environment could vary during episodes, not just between episodes as we have tested here. This is something that a GNN or Transformer-based policy such as RogueNet should also be robust to in principle.
The network structures underlying each instance of a Yawning Titan environment are constructed using the NetworkX package (Hagberg et al.,, 2008). Following Collyer et al., (2022), we generate these networks randomly using the Erdős-Renyi model with edge probability equal to $0.1$, with post-processing to make sure the graph is connected. Figure 1 shows examples of networks generated in this way for 10, 20 and 40 nodes, which are the network sizes we used for our experiments.
4.2 Termination
Both Andrew et al., (2022) and Collyer et al., (2022) consider environment configurations in which episodes terminate either when the red agent compromises a high-value node or when the episode reaches a maximum length. In our case, we fix the episode length at 100 environment steps, with no other termination conditions. This allows us to focus solely on the agent’s ability to generalise across different network topologies, without confounding factors from varying episode lengths.
4.3 Red Agent
In cyber defence simulators, the behaviour of the attacking red agent is typically generated in a rules-based fashion, either with some randomisation or entirely deterministically (Standen et al., 2022b, ; Andrew et al.,, 2022; Standen et al., 2021a, ; Standen et al., 2022a, ). The red agent is designed to spread through a given network and compromise nodes, while the blue agent is tasked with preventing this from happening.
For our environment, the red agent has two actions, Basic Attack and Zero Day, both of which target a single node per timestep:
  • Basic Attack targets a node and, if successful, allows the red agent to compromise it. Success depends on the ‘skill’ level of the red agent, defined as $s\in[0,1]$, and on the vulnerability score $v\in[0,1]$ of the targeted node. Together, skill level and vulnerability define the ‘attack strength’ $a$ as $a=\frac{100s^{2}}{s+(1-v)}$. The targeted node is compromised if $a>t$, where $t\sim\mathcal{N}(0,100)$, otherwise it fails.
  • Zero Day attacks allow the red agent to compromise a targeted node regardless of its own skill level or the vulnerability of the node. This gives the red agent the chance to make progress even when facing a strong defensive blue agent. However, so the basic attack is not redundant, this only becomes available once every three environment steps.
At each environment step, the red agent randomly chooses a node to act on. If the Zero Day attack action is available, this will be prioritised and used. Otherwise, the agent must resort to the Basic Attack.
Since the focus of our experiments was on varying network topologies, we decided to employ a fixed configuration for the red agent’s behaviour. However, the configuration and diversity of the behaviour of the red agent has a large impact on the difficulty of the task the blue agent is faced with, and the policy it will end up learning. For robustness, training a blue agent for real-world deployment would likely involve exposing it to a diverse range of red agent behaviour and configurations, including perhaps an intelligent attacker trained in opposition to the defender (Janisch et al., 2023a, ; Hammar and Stadler,, 2020; Microsoft Defender Research Team,, 2021).
4.4 Observations
The blue agent has access to the compromised status (a Boolean value) and the vulnerability score for every node, with no other information provided. For the agents trained on the standard Gym interface for Yawning Titan, the observation is a vector of these features for all nodes concatenated together. For the Entity Gym interface, the observation is given as a list, where each element of the list is the feature vector for the corresponding node. If there were different node types, the features would be split into separate lists for each node type.
4.5 Actions
The blue agent may only choose to act on one node per environment step. Following Collyer et al., (2022), we give the blue agent access to two actions: Reduce Vulnerability and Restore Node. In the standard OpenAI Gym configuration, the action space consists of a combination of each action with each node in the network, amounting to $2n$ actions if $n$ is the number of nodes.
  • Reduce Vulnerability allows the blue agent to reduce the vulnerability $v$ of a specific node by a fixed amount. The default amount is $0.2$, with a lower bound on the vulnerability $v$ placed at $0.01$, such that it cannot be reduced further. This is a proactive or preventative action, through which the blue agent can try to anticipate the red agent behaviour and strengthen nodes that are likely to be easily compromised.
  • Restore Node allows the blue agent to restore a node to its original state. This means that if the node has been compromised, the compromised status will be removed. While, if the vulnerability of the node has changed, it will be reset to its original value. This is a reactive action, through which the blue agent might respond to any inroads the attacker has made into the network.
4.6 Rewards
For the reward provided to the blue agent at each timestep, we again followed Collyer et al., (2022) and defined it as the proportion of nodes not compromised at timestep $t$:
$R_{t}=\frac{N-N^{c}_{t}}{N}$
where $R_{t}$ is the reward at timestep $t$, $N$ is the total number of nodes in the environment, and $N^{c}_{t}$ is the number of compromised nodes at timestep $t$. Defining rewards in this way induces a simple positive reward structure, which encourages the blue agent to minimise the number of compromised nodes. Collyer et al., (2022) also include a large negative reward of $-100$ for a high-value node being compromised, and a large positive reward of $100$ for an episode reaching its maximum number of steps. We did not make use of these additional rewards in our experiments, since our environment has fixed-length episodes and does not feature high-value nodes. With fixed episodes of length 100 under this reward structure, the maximum reward is 100, with a maximum per-timestep reward of 1 if no nodes are compromised.
5 Entity-based adaptation
In order to apply entity-based RL to Yawning Titan, we first adapted its underlying simulator to the Entity Gym interface. In this section we will provide a description of how this can be achieved.
5.1 Entity Types
In the Yawning Titan environment, it is possible to make the distinction between different types of nodes. Whilst all nodes behave in fundamentally the same way, there may be some nodes that have a special designation. These are the entry, high-value and deceptive nodes. For our experiments we used a single randomly-chosen entry node, with all others being generic. For this reason, we have chosen to use a single generic nodeEntity Type in the construction of the entity-based environment. In the RogueNet architecture (Winter et al.,, 2021), this means that a single shared entity encoder is used for all node embeddings during training.
In more realistic simulators or real networks, it may be beneficial to encode different types of network nodes using different Entity Type labels and therefore different encoders. This would allow the policy network to be more expressive and would provide a defensive agent with prior knowledge of the fundamental differences between nodes. It would also be impossible to encode all types of nodes with the same encoder if the features or data available to the agent on each node type differ. For example, in the CybORG environment (Standen et al., 2021b, ; Standen et al., 2022a, ), one might imagine designating separate entity types for the User, Enterprise Server and Operational Server nodes.
5.2 Entity Features
The blue agent receives observations consisting of a dictionary with keys for each entity type, each with a potentially variable-length list of entity features. For each node, we chose to include the vulnerability score and compromised status.
5.3 Action Space
In Yawning Titan, the blue agent may only take action to intervene on one node per time-step. The Entity-Gym framework offers three possible types of action, as outlined in Section 3.1. Using the entity-level CategoricalActionSpace means that a candidate action is chosen for every entity of a particular type at every environment step. In the case of games containing a large number of simultaneously acting entities like Starcraft, this is necessary to allow the agent to act on any subset of them simultaneously, in a tractable way. In our setting this would mean that the blue agent could direct any subset of nodes simultaneously to either reset themselves or reduce their vulnerability (with any combination of these actions for the selected nodes). This option may be desirable in real network security applications, but it disrupts the balance of the Yawning Titan game, making it significantly easier for the blue agent. It is likely this would be a similar issue in other autonomous cyber defence environments. This means it is necessary to create a two-stage, structured action space using a combination of a SelectEntityActionSpace and either a CategoricalActionSpace or a GlobalCategoricalActionSpace.
For selecting nodes to act on, we use a SelectEntityActionSpace, which requires defined actor and actee entity types. In order to use this action type, we create a dummy ‘Defender’ entity with fixed features to serve as an actor, which is able to act on all nodes in the network.
For the node-specific actions of Reduce Vulnerability and Restore Node, these can be implemented using either the ‘global’ categorical action space or the ‘entity-level’ categorical action space. For a GlobalCategoricalActionSpace, the RogueNet architecture creates a dummy global entity with random features, with information from the nodes of the network shared with its embedding vector through the attention mechanism in the Transformer layers. This single, compressed embedding is then provided as the input for the action head. Combined with the SelectEntityActionSpace, this means the agent chooses an ‘action type’ and then a node on which to execute it. Under the entity-level CategoricalActionSpace, all embeddings of the nodes are passed into the action head simultaneously in a batch, with a separate action selected for every node independently, based on each node’s respective embedding. Conceptually, combining this with a SelectEntityActionSpace would be like choosing a node to execute the action that it has selected, rather than choosing the action centrally. In our case, we chose to use the global action space, as this should offer marginally more efficient inference due to fewer actions being calculated. Incidentally, the underlying mechanism is very similar to the approach outlined in Janisch et al., 2023b , where environments are represented as a graph, and an extra ‘global node’ is instantiated to gather global information over a series of GNN message-passing steps. This global node is then used to choose the ‘action identifier’, or action type, before proceeding in an autoregressive fashion.
Wolk et al., (2022) and Hammar and Stadler, (2020) decompose the action space in a similar way, where a distinction is made between selecting a node and selecting an action type. Janisch et al., 2023b generalise this to selecting multiple nodes in the SysAdmin game, where each node is selected in sequence dependent on the last. In keeping with the architecture of RogueNet (Winter et al.,, 2021), we treat each component of the policy independently, similarly to Wolk et al., (2022). In other words, both the action type and node are chosen independently by different action heads, with the choices combined to execute the action. This is also similar to the decomposition chosen in Vinyals et al., (2017), and means the policy is simpler to learn in comparison to conditional action sequences. In more complex cyber environments, it might be beneficial to incorporate conditional action selection sequences, like the one used in AlphaStar (Vinyals et al.,, 2019; Choi,, 2020) or following the more general framework of conditional action trees proposed in Bamford and Ovalle, (2021), but we leave this for future research.
6 Experiments
In this section, we outline the methodology and present results from the experiments we conducted to compare a baseline MLP-based Proximal Policy Optimization (PPO) agent (Schulman et al.,, 2017) with an entity-based RogueNet (Winter et al., 2023a, ) policy, as well as experiments exploring the zero-shot generalisation abilities of entity-based agents in environments with network topologies not seen during training. We drew the baseline PPO agent implementation from Stable Baselines 3 (Raffin et al.,, 2021), which is a library that collects stable and performant implementations of popular deep RL algorithms, and is generally reported to have the most consistent performance among open-source PPO implementations (Huang et al.,, 2024; Huang et al., 2022a, ). We trained the RogueNet agents using the companion Entity Neural Network Trainer package (Winter et al., 2023c, ) which uses a simple implementation of PPO derived from the CleanRL library (Huang et al., 2022b, ).
Figure 2: Train-time episodic rewards evaluated on the 10, 20, and 40 node networks, respectively. The four agents compared in each evaluation are: baseline PPO agent on the static (sb3_static_[nodes]) and random (sb3_random_[nodes]) network environments, and the entity neural network agent on the same static (sb3_random_[nodes]) and random (‘Entity_random_[nodes]’) environments. Rewards are averaged as the mean over three different random seeds, and shaded error bands are constructed between the maximum and minimum of the three runs.
Figure 3: Box-plots of episodic rewards generated from the evaluations of three entity-based agents trained on different network sizes, over 1,000 test episodes. Subfigure (a) shows the evaluations of three entity-based agents trained on 10, 20 and 40-node networks respectively, and evaluated at test-time on random 10-node networks (eval_rand_k_on_10 with $\mathbf{k\in\{10,20,40\}}$). Subfigure (b) shows the same agents evaluated on 20-node networks (eval_rand_k_on_20 with $\mathbf{k\in\{10,20,40\}}$). Subfigure (c) shows the same agents evaluated on 40 node networks (eval_rand_k_on_40 with $\mathbf{k\in\{10,20,40\}}$).
mean std min 25% 50% 75% max eval_rand_10_on_10 93.840246 1.522998 91.000046 93.000046 93.500053 94.000046 98.900002 eval_rand_20_on_10 93.764247 1.229934 91.900063 93.100052 93.500053 94.000046 98.900002 eval_rand_40_on_10 92.848245 7.654986 16.100031 93.000053 93.500046 94.000046 98.900002 eval_rand_10_on_20 96.278664 0.547026 95.349945 95.999968 96.199966 96.399963 98.949997 eval_rand_20_on_20 96.232763 0.414635 95.099945 95.999960 96.199959 96.399956 98.949997 eval_rand_40_on_20 96.254164 0.453860 95.249954 95.999954 96.199974 96.399963 98.849991 eval_rand_10_on_40 97.591049 0.187861 97.199936 97.474945 97.574944 97.674957 98.974998 eval_rand_20_on_40 95.498728 0.332559 94.874931 95.274940 95.449917 95.624941 97.199944 eval_rand_40_on_40 97.601449 0.161840 97.224937 97.499947 97.599945 97.699951 98.875000
Table 1: Summary statistics for zero-shot episodic evaluation of trained entity-based agents on 1000 random networks of varying sizes. We evaluated 3 policies, trained on network sizes of 10, 20 and 40 respectively, on random networks of 10, 20, and 40 nodes, to assess their generalisation performance. In this table eval_rand_n_on_k refers to the entity-based policy trained on random n-node networks, evaluated on 1000 k-node networks.
6.1 Training
All of our experiments employed the Yawning Titan configuration described in Section 4, with variable randomised network topologies of 10, 20 and 40 nodes generated via the Erdős-Renyi method. We trained each agent on 10,000,000 total environment steps, which amounts to 10,000 episodes with a fixed length of 100 timesteps per episode. During training, we evaluated each policy every 10,000 environment steps and logged the resulting episodic reward to produce training graphs. Exchangeability. As suggested in Mern et al., (2020, 2019), one advantage of the entity-based approach and using a Transformer policy is ‘object exchangeability’. That is, the robustness of Transformers and GNNs to different permutations of objects in the observation vector or action space without the effectiveness of the policy being impacted. To demonstrate this property, we consider two different training settings, which we refer to as Static and Random. In the Static setting, a single network is randomly initialised at the start of training, and remains the same for every episode; the only thing that varies is the identity of the entry node. In the Random setting, an entirely new network is initialised for every episode. The aim of this setup is to evaluate policy robustness across inhomogeneous networks in the training phase, and is similar to the method used by Hong et al., (2022) in ‘Multitask Inhomogeneous Reinforcement Learning’ for robotic control.
In total, we trained policies from 3 random seeds for each combination of policy type and environment setting. Considering all the combinations of agents, network sizes and ‘Random’ or ‘Static’ environments, we have trained 36 policies:
  • sb3_static_$n$_node_network ×$\times 3$: Three different MLP policies trained using the Stable Baselines 3 PPO implementation, on static networks of 10, 20, and 40 nodes respectively. Each policy was initialised and trained using a different random seed.
  • sb3_random_$n$_node_network ×$\times 3$: Three different MLP policies trained using the Stable Baselines 3 PPO implementation, on $n$-node networks randomly generated with each episode reset for $n\in(10,20,40)$ (the Random) environment). Each policy was initialised and trained using a different random seed.
  • EntityYT_static_$n$_node_network ×$\times 3$: Three different Entity Neural Network RogueNet policies, trained using the entity-based version of Yawning Titan, on static networks of 10, 20, and 40 nodes respectively. Each policy was initialised and trained using a different random seed.
  • EntityYT_random_$n$_node_network ×$\times 3$: Three different Entity Neural Network RogueNet policies, trained using the entity-based version of Yawning Titan, on $n$-node networks randomly generated with each episode reset for $n\in(10,20,40)$ (the Random) environment). Each policy was initialised and trained using a different random seed.
Figure 2(a) shows the reward trajectories of all four types of policies trained on 10-node networks. Both entity-based agents (‘Entity_random_10’ and ‘Entity_static_10’) show similar reward trajectories that converge rapidly and plateau higher than their counterpart Stable Baselines MLP-based policies, close to the maximum possible reward of 100. The Stable Baselines PPO agents trained on the single static network (‘sb3_static_10’) converged quickly as well, but to a lower level of asymptotic rewards than the entity-based agents. In line with expectations, the baseline PPO agents trained on a series of random networks that vary across episodes (‘sb3_random_10’) struggled to learn a good policy and displayed high variance in their performance. This supports the notion that a Transformer entity-based policy is more robust to exchangeable representations, and performs no worse than when trained on a static environment, whereas a fixed MLP-based policy struggles to learn an effective policy when the object that each element of the agent’s observation vector represents changes with each episode (Mern et al.,, 2020, 2019; Zhao et al.,, 2022).
Figures 2(b) and (c) show similar results on 20-node and 40-node networks respectively. Here, the difference between the episodic reward of the two entity-based policies and the statically trained Stable Baselines PPO agent is less significant. This may be, in part, due to the nature of the reward function defined as the proportion of non-compromised nodes at time $t$. Since the skill and behaviour of the red agent are unchanged, this naturally means that a lower proportion of nodes will be compromised on a given timestep, and thus rewards for the blue agent will be higher in larger environments. Nevertheless, the Stable Baselines agent trained on randomly generated networks still does comparatively poorly, and struggles to learn a policy that is invariant to different network topologies.
6.2 Generalisation to unseen topologies
The second set of experiments we conducted was aimed at evaluating how well entity-based agents generalise at test-time to network sizes not seen during training. As the baseline MLP policies do not support dynamically-sized inputs that are different to the ones used at train-time, they cannot be evaluated on varying network topologies at test-time.
To assess the generalisation ability of entity-based policies across different network sizes, we took three policies trained in the Random training regime, one for each network size of 10, 20 and 40. We then evaluated each of the policies on random networks for each of the network sizes. This means that we evaluated a policy trained on random 10-node networks in the same context of random 10-node networks but also on 20 and 40-node networks (‘out-of-training’ context). All the evaluations were carried out over $1000$ episodes, and so $1000$ different random networks.
Figure 3 displays the distribution of episodic rewards over 1,000 test episodes in the form of box plots, obtained from the evaluations of the three RogueNet agents trained on different network sizes. Subfigure 3(a) presents evaluations of agents trained on 10, 20 and 40-node networks respectively, and evaluated on random 10-node networks at test-time (eval_rand_k_on_10 with $\mathbf{k\in\{10,20,40\}}$). Subfigures 3(b) and 3(c) presents the corresponding evaluations on 20-node networks, 3(b), and 40-node networks, 3(c). Summary statistics of the same experiments are presented in Table 1.
Overall, we can observe a negligible difference in the performance of the three entity-based agents across the different network sizes when evaluated at test-time. For example, the agent trained on 40-node networks performs just as well when evaluated on 10-node networks as the agent that was trained natively on 10-node networks. In fact, it shows slightly lower variance in episodic reward than the agent trained on 10 nodes, with $0.87$ standard deviation against the $1.31$ standard deviation of the native 10-node agent. The only exception is the agent trained on 20-node networks and evaluated on 40-node networks, which shows slightly lower mean reward compared to the other two agents. As can be seen from the summary statistics results in Table 1, this case presents higher standard deviation in the rewards distribution (0.332) compared to the other two cases (0.187 and 0.161). These results suggest an ability for an entity-based agent to learn a general policy that performs consistently on any network size. The generality of these results would need to be verified in more complex environment settings or simulators, and different varieties of network topologies.
7 Discussion
In this paper, we have demonstrated the basic functionality and the strong generalisation capabilities of an entity-based RL method through the use of the Entity Gym framework for the Yawning Titan autonomous cyber defence environment. We suggest that other, similar or more complex cyber defence simulators could be readily adapted in the same way, due to the fundamental node-based structure of the observation and action spaces. In fact, when designing a simulator, it is likely to be much more straightforward to use an interface such as Entity Gym than devising rules for how a fixed-input agent can interact through the standard Gymnasium interface. One drawback is the lack of out-of-the-box implementations of algorithms designed to work with entity-based environments, and tuned for optimal performance with complex Transformer or GNN policies. Indeed, as can be seen from the training plots, the RogueNet policy displays occasional instability during training in the environment setting we used. However we believe this is outweighed by the potential for training and generalisation across variable network topologies, and can be rectified with more development and testing. We suggest that future cyber defence environments be designed using an ‘entity-based’ philosophy, and constructed using an interface akin to Entity Gym.
8 Future work
8.1 Network Information
If the network structure is known to the defender, it could be beneficial to provide this information to the blue agent. This could potentially be in the form of incorporating structural bias into the policy network by using a GNN or Graph Attention Network (Palmer et al.,, 2023). In the RogueNet architecture, introducing attention masking over node connections to reflect the structure of the network would be a relatively straightforward change.
Alternatively, as in continuous control (Kurin et al.,, 2021), if it is detrimental to restrict message-passing to neighbouring nodes, it is also possible to include global features that contain network connectivity information. Similarly to Hong et al., (2022), one might use a full Transformer policy, but with a graph-based positional encoding scheme.
8.2 Environment complexity and intra-episode variation
We have demonstrated the ability of the entity-based framework and a Transformer-based policy to train on varying network topologies with constant nodes, as well as to generalise zero-shot to network sizes not seen during training. It is also straightforward to train a policy on varying numbers of nodes during training, where the agent sees a different-sized network with every episode. The agent performs well and also converges to a high reward, although we don’t include this here since the reward function is dependent in-part on the size of the network. In the future, one might expect this kind of variety to be important for training robust agents.
Another important benefit of the Entity Gym, and entity-based or object-oriented RL more generally, is the support for natural variation in the population of entities within an episode, not just between episodes. In a deployment on a real network, the number of nodes observed by a defensive agent may vary over time. This could be due to devices dropping in and out of wireless connection, being manually connected or disconnected by administrators, or communication being blocked by an adversary. In the Yawning Titan environment, it is possible to enable the use of deceptive nodes by the blue agent, which would allow the agent to add new nodes to the environment mid-episode.
9 Acknowledgments
The authors would like to acknowledge helpful feedback and comments from Liz Bates and Ed Chapman. This work was supported by the Defence and Security Programme at The Alan Turing Institute, funded by the Government Communications Headquarters (GCHQ). Alberto Caron was funded by the Defence Science and Technology Laboratory (DSTL) which is an executive agency of the UK Ministry of Defence providing world class expertise and delivering cutting-edge science and technology for the benefit of the nation and allies.